OpenStudy (anonymous):

22. A ball is thrown into the air with an upward velocity of 24 ft/s. Its height (h) in feet after t seconds is given by the function h = –16t2 + 24t + 5. How long does it take the ball to reach its maximum height? What is the ball’s maximum height? Round to the nearest hundredth, if necessary. (1 point) a. 1.5 s; 5 ft. b. 0.75 s; 32 ft. c. 0.75 s; 11 ft. d. 0.75 s; 14 ft.

OpenStudy (anonymous):

\[h = –16t^2 + 24t + 5\] is the equation for a parabola. the vertex of the parabola always has the first coordinate \(-\frac{b}{2a}\) which in your case is \(-\frac{24}{2\times (-16)}=\frac{3}{4}\]

OpenStudy (anonymous):

i means which in your case is \(-\frac{24}{2\times (-16)}=\frac{3}{4}\)

OpenStudy (anonymous):

replace \(t\) by \(\frac{3}{4}=.75\) to find the height